# Using Fiber Optic Bundles to Miniaturize Vision-Based Tactile Sensors

### v2

@@ V-B Classification Results
To train the classification model, we first prepared the dataset—the raw RGB frames are cropped and downscaled to 224 × 224 pixels. To create the training, validation, and test sets we dedicated from each video frame sequence the first 80% of the frames for training and the last 20% of the frames for validation and testing. From the initial 80% of the video sequence, we create the training set by sampling from overlapping sequences of 16 frames with stride 5. Validation and test sets were created from the remaining 20% of the video sequence by sampling non-overlapping sequences of 16 frames, where the first 10% is for validation and the remainder for test.
@@
To perform classification, we fine-tuned transformer-based video masked autoencoders (VideoMAE) [51]. Fine-tuning was performed by training of a linear model on base model representations. The model takes as input the 16 frames sequences, and outputs a scalar that represents the hardness class. We implemented our models with PyTorch and the large instance of a pre-trained VideoMAE model openly available on HuggingFace. We used Adam optimizer with learning rate 0.005 and batch size 8, and a cross entropy loss for the classification task. The results are shown in the full confusion matrix (Fig. 9).
@@
After evaluating the trained classification model on the test set, the model was capable of reaching 97.8% accuracy over 9 hardness classes. Based on the results, we see that the trained model performs well in classifying elastomer hardness values that correspond to the range from healthy to unhealthy tissue.
@@ VI-A Classification of cancerous lesions in phantom prostate
We desired to test if DIGIT Pinki could make clinically relevant distinctions between cancerous lesions and healthy tissue in a prostate phantom. Six phantom prostates were first prepared by creating a mold from a patient prostate model, open-sourced by the Johns Hopkins Urology Robotics Program, NCI 5R01CA247959. We selected the prostate model, as this was the closest size corresponding to the average volume over four radical prostatectomies that we witnessed for procedures at Stanford Hospital. We only created a mold from the posterior half, which is the portion that would be palpated in a digital rectal exam. The mold is 3D-printed using a FormLabs3 printer with Rigid 1K resin and prepared with XTC-3D brush-on finish.
We cast the healthy portion of the posterior prostate using Ecoflex 0050 (50 Shore 00), which was chosen for having a similar elastic modulus to healthy prostate from patients [50], producing a 0.5A hardness value with a handheld durometer. We colored the Ecoflex 0050 silicone pink using 0.5% Silc-Pig red and 0.25% Silc-Pig white by weight. During the curing process, we then embedded harder spherical nodules of , , and inside the phantom at a depth of to mimic a cancerous lesion. The mixtures for the spherical nodules were mixed with the same ratios detailed in the Appendix C. The phantom prostates are fully cured at room temperature after 4 hours.
We chose to embed five different nodule hardness values inside the phantom prostates (45A, 25A, 16A, 11A, and 6.5A) to capture a range of lesion instances. As in the previously described experiment, we manually verified the hardness values of the nodules by taking the average of five measurements with a durometer from a flat silicone piece cast in the same batch as the nodule. As for the choice of hardness value for the nodules, a previous study on DRE-detectable nodules used hardness values of 23A, 27A, and 31A in simulated tissue [52], but in consultation with clinicians at Stanford Hospital who regularly perform the DRE on patients, we learned that prostate hardness values for healthy and unhealthy tissues vary widely from patient to patient. Therefore, we desired to make phantoms that reflected a wider range of possible lesion hardness values. The verisimilitude of the final phantom prostates was qualitatively assessed with clinician feedback. An example of one of the prostate phantoms is pictured in Fig. 10.
@@
The classifier model and parameters were set up as described in the previous experiment in Section V. We trained the classifier on the phantom palpation dataset, which included all embedded nodule sizes (4, 8, and diameter), the six detected hardness classes (where 0.5A is the baseline healthy tissue), and palpations with a gloved sensor and an ungloved sensor. As shown in Fig. 11, we achieved a perfect accuracy score on the classification task, suggesting good performance in hardness classification on phantom prostates. We note that the classification worked well even when the sensor was gloved. Furthermore, the classification task is able to achieve respectable performance even given the different simulated lesion sizes. A human clinician might have trouble feeling the smallest nodule during a real exam [52].
@@ VI-B Classification of cancerous lesion in ex vivo prostate tissue
We measured perfect accuracy on the test set, with Fig. 13 showing the confusion matrix. Because we were only able to sample one real prostate specimen, the classification result here is of limited statistical significance. However, the result suggests that the tactile sensor is able to distinguish between the two sampled points with different stiffness values and supports the suitability and clinical relevance of DIGIT Pinki on real tissue.
@@ VII Discussion and Future Work
In this work, we have demonstrated a new tactile sensor design that enables miniaturization beyond that achieved by most vision-based tactile sensors. The work is motivated by a desire to create fingertips with dimensions no larger than those of human fingertips, while maintaining the spatial and force resolution that have made vision-based tactile sensing popular. The resulting sensor is suitable for manipulation and exploration tasks where space is constrained, including the palpation of human organs.
VII-A Future Design Directions
Ample further opportunity exists for tuning a fiber-based approach specifically to tactile sensing. One direction includes researching different fiber and microlens architectures, taking advantage of the optical customization that fibers offer. An immediate extension is to test the hyperfisheye lens design introduced in this work and to increase the fiber core count accordingly, either through purchase or by manually heating, stretching, and combining discrete fibers.
However, one does not necessarily need to use a hyperfisheye lens to achieve a short working distance at the distal end. An interesting design direction is to replace the lens with a microlens array, producing an optical configuration akin to the compound eyes commonly found in arthropods.
@@ VII-B Additional Applications
Within the realm of medical applications, we would like to investigate other patient examinations that require understanding the tactile properties of tissue. For example, cervical tissue stiffness is an important indicator for preterm births and this sensor could be used to palpate cervical tissue to characterize stiffness. Similarly, during an open surgery of the pancreas, the surgeon will often palpate the pancreas because texture is an important predictive factor of post-operative complications. A robotic counterpart could provide repeatable quantitative measurements of texture by using a sensor like DIGIT Pinki.
@@ VIII Conclusion
In this work, fundamental design considerations have been investigated for an fiber-based tactile sensor, including fiber illumination and fiber spatial resolution. The sensor image-to-force regression model shows good accuracy for a given indentation, illustrating a potential for use in closed-loop manipulation tasks in constrained environments. The image-to-hardness classification models demonstrate the potential for a miniature high-resolution tactile sensor in tissue palpation through internal orifices for patient examinations, such as for the cervix or the prostate. We open-source the design and manufacturing process of DIGIT Pinki at https://github.com/facebookresearch/digit-design.
@@ Appendix A Complete CAD Assembly
We provide the full CAD model for the DIGIT Pinki setup in Fig. 15.
Appendix B Sensor Construction Manufacturing Guide
@@
Appendix C Silicone Preparation for Hardness Experiments
We include in Table III the mixing ratios used to achieve the silicone samples of different hardness values for this paper. After selected the desired range of hardness values, we specifically chose to use Smooth-Sil 945 and Ecoflex 0050 for manufacturing ease; they both have similar cure times and one-to-one mixing ratios between Parts A and B.
@@ Appendix C Silicone Preparation for Hardness Experiments
Hardness Value |
Smooth-On Ecoflex 0050 () |
Smooth-On Smooth-Sil 945 () |
50 (Shore 00) |
16 |
0 |
1.5 (Shore A) |
14 |
2 |
5 (Shore A) |
12 |
4 |
6.5 (Shore A) |
11.5 |
4.5 |
11 (Shore A) |
10 |
6 |
16 (Shore A) |
8 |
8 |
20 (Shore A) |
7 |
9 |
25 (Shore A) |
6 |
10 |
30 (Shore A) |
5 |
11 |
35 (Shore A) |
4 |
12 |
42 (Shore A) |
2 |
14 |
45 (Shore A) |
0 |
16 |

### v3

@@ V-B Classification Results
To train the classification model, we first prepared the dataset—the raw RGB frames are cropped and downscaled to 224 × 224 pixels. To create the training, validation, and test sets we dedicated from each video frame sequence the first 80% of the frames for training and the last 20% of the frames for validation and testing. From the initial 80% of the video sequence, the training set was created by sampling from overlapping sequences of 16 frames with stride 5. Validation and test sets were created from the remaining 20% of the video sequence by sampling non-overlapping sequences of 16 frames, where the first 10% is for validation and the remainder for test.
@@
To perform classification, we fine-tuned transformer-based video masked autoencoders (VideoMAE) [53]. Fine-tuning was performed by training of a linear model on base model representations. The model takes as input the 16 frames sequences, and outputs a scalar that represents the hardness class. The models were implemented with PyTorch and the large instance of a pre-trained VideoMAE model openly available on HuggingFace. We used Adam optimizer with learning rate 0.005 and batch size 8, and a cross entropy loss for the classification task. The results are shown in the full confusion matrix (Fig. 8).
@@
After evaluating the trained classification model on the test set, the model was capable of reaching 98% accuracy over 11 hardness classes. Based on the results, the trained model performs well in classifying elastomer hardness values that correspond to the range from healthy to unhealthy tissue.
V-C Analysis of Contact Surface Properties and Image Sequence Length for Hardness Classification
To further study the learning ability of the VideoMAE model, rather than using the tactile images, delta-frames were instead used as input, computed as the difference between the current frame and the next frame (i.e. ). Using delta-frames ensured that the model was learning mostly from the evolution of the contact patch and not spurious artifacts. Training on the 16-frame sequence of delta-frames yielded an accuracy of 94.4% over 11 hardness classes, comparable in performance to that from training with the tactile images. This result gave confidence that the model was learning to classify hardness based on the contact patch, rather than obscure artifacts not grounded in physical properties of the contact.
We note, however, that perceived material stiffness is a function of both material hardness and geometry, and yet the classification results were quite accurate despite the different diameter geometries in the sample dataset. Although the evolution of the contact patch was important, it was likely not the only discriminant to the machine learning model. To show this, we highlight an adversarial case: a sample with a large diameter bump made of a stiff material and a sample with a small diameter bump made of pliable material could have the same contact patch size when indented. This scenario was represented in the collected silicone hardness dataset, and was modeled using COMSOL Multiphysics, a finite element analyzer. Because the scenario is radially symmetric, only half of the indenter and sample were modeled for computation speed, with the model parameters given in Appendix G. As shown in Fig. 9, although the contact patch radius was the same, the different bump diameters lead to different surface pressure distributions when indented with a load. Even in adversarial cases where the sensor may image a similar contact patch, there was sufficient resolution to capture other physical properties of the contact, explaining in part how the model discriminates hardness classes accurately despite different sample geometries.
We also note that in the ideal case, indenters should be matched in mechanical impedance with their samples for optimal sensitivity [54, 55]. In this work, the Shore 10A indenter material (around Shore 00-60) was 10 degrees harder than the softest samples in the dataset (Shore 00-50). There was therefore a reduction in sensor sensitivity because DIGIT Pinki used a stiffer-than-optimal material for the softest samples. However, the material mismatch was not severe enough to significantly degrade performance on soft samples. We provide more analysis in Appendix H to explore the tradeoff between indenter material and sensitivity.
A study was also conducted on the image sequence length. In this study, the model used the full 16-frame sequence of delta-frames as input, where the last frames in the sequence were masked by black frames for . This corresponded to effective image sequences of 16, 8, 4, 2, and 1 delta-frames respectively. The F1 score results are plotted in Fig. 10 and the accuracy scores in Appendix I. Based on the results, the VideoMAE model performed best with longer image sequences, suggesting that sequence length is important for task performance.
@@ VI-A Classification of cancerous lesions in phantom prostate
We desired to test if DIGIT Pinki could make clinically relevant distinctions between cancerous lesions and healthy tissue in a prostate phantom. Six phantom prostates were first prepared by creating a mold from a patient prostate model, open-sourced by the Johns Hopkins Urology Robotics Program, NCI 5R01CA247959. The 2 prostate model was selected, as this was the closest size corresponding to the average volume over four radical prostatectomies that were witnessed at Stanford Hospital. A mold was created only from the posterior half, which is the portion that would be palpated in a digital rectal exam. The mold is 3D-printed using a FormLabs3 printer with Rigid 1K resin and prepared with XTC-3D brush-on finish to ensure that the surface was smooth.
We cast the healthy portion of the posterior prostate using Ecoflex 0050 (50 Shore 00), which was chosen for having a similar elastic modulus to healthy prostate from patients [52], producing a 0.5A hardness value with a handheld durometer. We colored the Ecoflex 0050 silicone pink using 0.5% Silc-Pig red and 0.25% Silc-Pig white by weight. During the curing process, we then embedded harder spherical nodules of , , and inside the phantom at a depth of to mimic a cancerous lesion. The mixtures for the spherical nodules were mixed with the same ratios detailed in the Appendix E. The phantom prostates are fully cured at room temperature after 4 hours.
We chose to embed five different nodule hardness values inside the phantom prostates (45A, 25A, 16A, 11A, and 6.5A) to capture a range of lesion instances. As in the previously described experiment, we manually verified the hardness values of the nodules by taking the average of five measurements with a durometer from a flat silicone piece cast in the same batch as the nodule. As for the choice of hardness value for the nodules, a previous study on DRE-detectable nodules used hardness values of 23A, 27A, and 31A in simulated tissue [56], but in consultation with clinicians at Stanford Hospital who regularly perform the DRE on patients, it was apparent that prostate hardness values for healthy and unhealthy tissues vary widely from patient to patient. Therefore, it was desireable to make phantoms that reflected a wider range of possible lesion hardness values. The verisimilitude of the final phantom prostates was qualitatively assessed with clinician feedback. An example of one of the prostate phantoms is pictured in Fig. 11.
@@
The classifier model and parameters were set up as described in the previous experiment in Section V. We trained the classifier on the phantom palpation dataset, which included all embedded nodule sizes (4, 8, and diameter), the six detected hardness classes (where 0.5A is the baseline healthy tissue), and palpations with a gloved sensor and an ungloved sensor.
As shown in Fig. 12, we achieved a perfect (100%) accuracy score on the classification task, suggesting good performance in hardness classification on phantom prostates. We note that the classification worked well even when the sensor was gloved. Furthermore, the classification task is able to achieve respectable performance even given the different simulated lesion sizes. A human clinician might have trouble feeling the smallest nodule during a real exam [56].
For the phantom prostate data, we also repeated the studies on contact patch evolution and the image sequence length using the same delta-frame input preparation as described in Section V-C. Training on a 16-frame sequence of delta-frames yielded an accuracy of 97.97%, giving confidence that the model was learning from the physical properties of the contact with the phantom prostate.
Similarly, the effects of image sequence on performance were evaluated with sequences of masked frames where , corresponding to 16, 8, 4, 2, and 1 delta-frame sequences. The F1 scores and accuracy scores are also included in Appendix I. As with the silicone hardness samples, the model performed best on the phantom prostate samples with longer image sequences.
@@ VI-B Classification of cancerous lesion in ex vivo prostate tissue
We measured perfect accuracy on the test set, with Fig. 14 showing the confusion matrix. Despite the more complex geometry of the prostate, it was not entirely a surprise that DIGIT Pinki was able to perform well. Previous literature has shown that vision-based tactile sensors perform better at hardness estimation on smooth and simple rather than sharp geometries [55, 54]. Prostate glands are smooth bulk tissues with no sharp edges (the dark-colored blood and tissue on the surface can be visually confounding and may give the appearance of complicated geometries in photographs), so they deform predictably. Because we were only able to sample one real prostate specimen, the classification result here did not represent a full medical study. However, the result suggests that the tactile sensor is able to distinguish between the two sampled points with different stiffness values and supports the suitability and clinical relevance of DIGIT Pinki on real tissue.
@@ VII Discussion and Future Work
In this work, we have demonstrated a new tactile sensor design that enables miniaturization beyond that achieved by most vision-based tactile sensors. The work is motivated by a desire to create fingertips with dimensions no larger than those of human fingertips, while maintaining the spatial and force resolution that have made vision-based tactile sensing popular. The resulting sensor is suitable for manipulation and exploration tasks where space is constrained, including the palpation of human organs. The experimental results demonstrate that DIGIT Pinki can learn clinically-relevant information from a representative palpation task, and provides merit for learning the details of the contact patch as part of diagnostic procedure.
VII-A Limitations
The proposed design is not without limitations. A major limitation is cost; manufacturing a small hyperfisheye lens prototype in limited quantities is expensive, so only the tip of the sensor could be imaged. The cost, however, would decrease when produced in bulk. Furthermore, finished coherent imaging fiber bundles and optical imaging lens pieces are also expensive, though raw fiber spools may be purchased at a cheaper rate.
Another limitation of this design is that miniaturization generally sacrifices resolution. In this design, because the fiber cores are the pixels of an image, the quality and quantity of the fiber cores dictate the resulting resolution. Although DIGIT Pinki has enough spatial resolution to perform the palpation task accurately, it could have difficulty performing other manipulation tasks that require much higher resolution, such as very fine-grained texture classification. A broader exploration of how much resolution is needed for certain tasks is out of scope, but warrants further investigation.
Finally, fiber bending must be considered. The fibers themselves have a minimum bend radius (to ensure no breakage or light loss). It is reasonable to think that this sensor design as is would not suitable for high-impact forces or for settings where the fibers may experience sharp bends beyond the minimum bend radius, which generally scales inversely with cost. For context, in this work, the minimum bend radius was (but this was not optimized during the design process because clinicians do not bend their fingers during rectal examinations). It is relatively straightforward to buy affordable fiber bundles (a few U.S. dollars per 10 ft.) with minimum bend radii down to , but beyond that they could become costly.
This work also did not investigate fatigue or cycle testing with repeated bending of fibers, which we leave as future work. However, extensive mechanical fatigue testing of optical fibers has been conducted in the literature to millions of cycles [57], demonstrating better fatigue performance than copper wire common in prototyping [58]. Defects in the silica can degrade and propagate if the fibers become wet, however almost all commercially-available fibers (including those used in this work) are sheathed with a protective jacket to prevent degradation from moisture and abrasion.
VII-B Future Design Directions
Ample further opportunity exists for tuning a fiber-based approach specifically to tactile sensing. As noted earlier, it would be desireable to substitute the lens in this prototype with a hyperfisheye lens, assuming reasonable cost, to image the whole surface of the sensor. Details on the considerations for designing a hyperfisheye lens specifically for tactile sensing are noted in Appendix J. Another direction includes researching different fiber and microlens architectures, taking advantage of the optical customization that fibers offer. An immediate extension to increase the fiber core count, either through purchase or by manually heating, stretching, and combining discrete fibers.
However, one does not necessarily need to use a standard lens to achieve a short working distance at the distal end. An interesting design direction is to replace the lens with a microlens array, producing an optical configuration akin to the compound eyes commonly found in arthropods.
@@ VII-C Additional Applications
Within the realm of medical applications, we would like to investigate other patient examinations that require understanding the tactile properties of tissue. For example, cervical tissue stiffness is an important indicator for preterm births and this sensor could be used to palpate cervical tissue to characterize stiffness. Similarly, during an open surgery of the pancreas, the surgeon will often palpate the pancreas because texture is an important predictive factor of post-operative complications. A robotic counterpart could provide repeatable quantitative measurements of texture by using a sensor like DIGIT Pinki. One may modify DIGIT Pinki’s gel to be selectively transparent or opaque, for applications where one may need standard imaging in addition to tactile imaging in constrained settings.
@@
The approach of the DIGIT Pinki design to separate the sensing element from supporting circuitry could also be expanded upon for applications with multiple fingertips or even for robotic skins. A future dexterous robot hand with multiple sensorized fingertips, in lieu of using multiple traditional vision-based tactile sensors (requiring a camera, USB cable, and associated wiring management per finger), could instead have multiple passive gel tips combine to a single optical fiber bundle, camera, and USB connection downstream. A fully sensorized robot hand with tactile fingertips and a large-area skin (for the palm or back of the hand), could also be made possible with multiple optical fiber bundles. Although routing a fiber bundle may prove challenging, a fiber-based design could alleviate the wiring/packaging problem of many existing tactile glove efforts. To achieve the short focal distance, a DIGIT Pinki skin design may take advantage of a microlens approach.
@@ VIII Conclusion
In this work, fundamental design considerations have been investigated for an fiber-based tactile sensor, including fiber illumination and fiber spatial resolution. The sensor image-to-force regression model shows good accuracy for a given indentation, illustrating a potential for use in closed-loop manipulation tasks in constrained environments. The image-to-hardness classification models demonstrate the potential for a miniature high-resolution tactile sensor in tissue palpation through internal orifices for patient examinations, such as for the cervix or the prostate. We open-source the design and manufacturing process of DIGIT Pinki at https://github.com/facebookresearch/digit-design, and provide a video summary at https://youtu.be/JaLls5nyrX4.
@@ Appendix A Complete CAD Assembly
The full CAD model for the DIGIT Pinki setup is provided in Fig. 16. The CAD files are also open-sourced at https://github.com/facebookresearch/digit-design.
Appendix B Assembly Considerations and Modifications
As previously mentioned, the screw-on gel was made possibly because of the modularity of this fiber-based design, which allowed the circuitry to be entirely separated from the gel tip. This design was especially useful in a medical context because the gel tip could be easily replaced after each use. There was a concern that because the gel was not directly overmolded to the surface, the connection to the rest of the sensor system was not as secure. Depending on the tolerance of the 3D-printed parts, some cheaper printers were not able to print a fine enough thread, no such issues were noticed the final prototype. The following settings were used for all 3D-printed parts:
| Component | Specification |
| 3D Printer | Stratysys J750 |
| Layer Height | Default |
| Material | Vero Black Plus |
| Finish | Glossy |
In addition, once the prototype was assembled, the image focus could be adjusted in a number of ways. The mounts for the microscope eyepiece and lens are mounted directly to the optical breadboard with metric screws; however, a 3D-printed part was designed with slots instead of mounting holes such that the pieces could slightly slide forward or backward to account for any misalignment before being tightened to a final position on the optical breadboard base. This allowed the exact placement of each component in the imaging assembly to be manually adjusted, and these 3D-printed mount designs are open-sourced. The adjustable diopter could also be used to adjust focus, but only by a small amount because of the shorter travel, so we relied on lens placement for any greater adjustments. A ThorLabs aluminum alignment target was used to help with adjustments, but because the process was manual, slight variations in focus occurred.
To bring DIGIT Pinki to the Stanford Hospital for the prostatectomy observation, it was required that all instrumentation remain relatively sterile/clean-able and strictly handheld. The design was made portable by placing the imaging assembly and lighting in a small plastic bin that could carried in and covered with a sterile bag. A hole was punched through the container to feed the fibers through, a second hole was punched through to feed the USB cable. All potential blood-contacting surfaces were covered with nitrile glove material. Because of the tight quarters of the operating room, image data was collected using a USB camera phone app on an Android device rather than a computer, but the image resolution and sampling frequency rate was matched to that of the benchtop data collection setup.
Appendix C Sensor Construction Manufacturing Guide
@@
Appendix D Robot Indenter Data Collection Setup
To collect force data, several other vision-based tactile sensors used a ground truth sensor embedded under the tactile sensor prototype, so that the forces from the indenter probe onto the tactile sensor were measured. As a quasi-static system, a rigid frame transformation could be calculated between the tactile sensor and the ground truth sensor.
In this work, the ground truth sensor could have been placed under the DIGIT Pinki on the robot arm end effector. However the extruding fibers of DIGIT Pinki made it not so straightforward to mount a ground truth sensor underneath, leading to an alternative configuration that instead placed the ground truth sensor under the indenter probe. In this case, the ground truth sensor measured the forces imparted on the indenter probe by the DIGIT Pinki, which was translated into the probe using the Meca500 robot arm. The force balance equation was still . Because the axes of DIGIT Pinki and the ground truth sensor were aligned, and only translation was applied, the frame transformation is straightforward to calculate.
Appendix E Silicone Preparation for Hardness Experiments
We include in Table IV the mixing ratios used to achieve the silicone samples of different hardness values for this paper. After selecting the desired range of hardness values, we specifically chose to use Smooth-Sil 945 and Ecoflex 0050 for manufacturing ease: they both have similar cure times and one-to-one mixing ratios between Parts A and B.
@@ Appendix E Silicone Preparation for Hardness Experiments
| Hardness Value | Smooth-On Ecoflex 0050 () | Smooth-On Smooth-Sil 945 () |
| 50 (Shore 00) | 16 | 0 |
| 1.5 (Shore A) | 14 | 2 |
| 5 (Shore A) | 12 | 4 |
| 6.5 (Shore A) | 11.5 | 4.5 |
| 11 (Shore A) | 10 | 6 |
| 16 (Shore A) | 8 | 8 |
| 20 (Shore A) | 7 | 9 |
| 25 (Shore A) | 6 | 10 |
| 30 (Shore A) | 5 | 11 |
| 35 (Shore A) | 4 | 12 |
| 42 (Shore A) | 2 | 14 |
| 45 (Shore A) | 0 | 16 |
Appendix F Spatial Points For Study of Separated Locations
We provide here the separated train/val/test locations for the smaller 10k dataset study. Within the region centered on the sensor tip, the validation spatial points for the , , and Flat indenters were , , and respectively; for the test set points, they were , , and respectively. The training points were a sweep along the x-axis at increments of from to .
Appendix G COMSOL FEA Simulation Parameters
Figure 9 includes the results of the FEA plots which studied an adversarial case when the contact patch size is the same. The modeling was done in COMSOL Multiphysics 6.2. The exact geometries of sensor tip and the silicone samples were used. Because the scenario was radially symmetric, only one half of the indenter and sample was modeled to save computation time. The indenting tip was modeled with a Young’s modulus of , a Poisson’s ratio of 0.49, and a density of based on previously published values for Solaris [59]. The Ecoflex 00-50 sample was modeled with a Young’s modulus of , a Poisson’s ratio of 0.36, and a density of [60]. When indenting, the sensor tip applied a ramped load up to normally into the sample. The indenter tip contact boundary is meshed at 0.3 and the sample contact boundary is meshed at 0.15.
Appendix H Additional FEA Study on Tradeoff of Mismatched Indenter and Sample Hardnesses
As noted in Section V-C, the DIGIT Pinki material was not optimized for the softest tested samples, as there was an on the order of 10 degree difference in Shore hardness between the indenter (Shore 10A or about Shore 00-60) and the softest sample (Shore 00-50). This difference was not significant enough to degrade the performance on the softest samples in this work. However, previous research on tactile hardness estimation has shown that when there was a more significant mismatch between the indenter and sample hardness, the accuracy of the hardness estimation was reduced [54]. Previous research has also shown that stiffer gel construction can be more robust against abrasion [1], attractive for a critical application such as medicine. To better indicate the price of using a stiffer-than-optimal indenter material on soft samples, we report an additional FEA study. The study compared the deformation at the surface for presses of 0 to under three scenarios: (a) an Ecoflex 00-50 indenter, (b) the DIGIT Pinki (Shore 10A or Shore 00-60) indenter, (c) and a plausible SORTA-Clear 40 (Shore 40A or Shore 00-86) indenter all on a diameter Ecoflex 00-50 sample. Based on the results, shown in Fig. 17, the (a) impedance-matched scenario is the most sensitive; a light indentation of causes a surface displacement of about . To achieve the same displacement, the (b) current sensor setup must be pressed with , showing that the mismatch in material does lead to a decrease in sensitivity. In scenario (c) where the mismatch approached that of a 50x magnitude difference, there was practically no deformation for light forces. This result suggested that increasing palpation force could help compensate for small mismatches in material hardness on the order of 10 degrees, but not for larger mismatches.
Appendix I Additional Results From Studying Image Sequence Length
Figure 18 show the results for the delta-frame study on the importance of image-sequence length to hardness prediction success, where the delta-frames were computed as . This study reported the accuracy scores on silicone hardness data, and both the F1 score and accuracy scores for the phantom hardness data. The results indicated that longer image sequences were important for the model to accurately predict hardness.
Appendix J Hyperfisheye Lens Design Considerations for Miniature Tactile Sensors
There are some considerations that must be made to design a compact hyperfisheye lens specifically for tactile sensing, such as the theoretical field-of-view lens shown in Fig. 19. For a dome-shaped sensor made of bulk silicone, the lens must be modeled for a different index of refraction than air. Compared to commercially-available lenses intended for photography, lenses for tactile sensing are relatively unconcerned about optical distortions. Furthermore, intentionally allowing chromatic aberrations and removing anti-reflective coatings better captures reflection and scattering inside the elastomer. The design modeled here uses APEL 5014 plastic material and could be manufactured either by diamond turning or molding.
The MTF performance curves are modeled in Code V and illustrated in Fig. 20, relating contrast over spatial frequencies (resolution) at the wavelength. As a lens for a dome-shaped tactile sensor, the design optimizes for high MTF across the field of view. As seen in the MTF curves, the imaging resolution for an indentation at the sides will be slightly worse than the performance for an indentation at the top. For example, the worst resolving performance (about 170 cycles/) occurs at the maximal diffraction angle of , corresponding to an indentation at the very edge of the field of view. However, when compared with the resolving performance for an indentation at the top of the sensor (about 200 cycles/), this is only about a 15% reduction in resolution.
Due to resource constraints, we were unable to manufacture the hyperfisheye lens design in time for testing. Although the custom lens design presented was not integrated in the sensor system, the MTF analysis presented should accurately reflect the real-world MTF when manufacturing tolerances are performed correctly. While the spatial resolution result reported in Section IV is with the lens, the same resolution with a hyperfisheye lens could be attained by increasing the fiber core count accordingly.